OpenStudy (adi3):

Will Medal Please help solve for x: x-1/1+2 = 2/x

OpenStudy (adi3):

@IrishBoy123

OpenStudy (irishboy123):

can you draw it, do not understand what you are solving...

OpenStudy (adi3):

solve for x: \[\frac{ x-1 }{ x+2} = \frac{ 2 }{ x }\]

OpenStudy (irishboy123):

multiply both sides by \((x+2)x\) [or cross multiply, as it is called, though it is the same thing...]

OpenStudy (adi3):

i am going with cross, x(x-1) = 2(x+2)

OpenStudy (irishboy123):

great

OpenStudy (adi3):

ok, then

OpenStudy (irishboy123):

multiply out and combine terms to get a quadratic

OpenStudy (adi3):

x^2 - 1x = 2x+4

OpenStudy (adi3):

then?

OpenStudy (irishboy123):

then \(x^2 - 3x - 4 = 0\) and now factor

OpenStudy (adi3):

= (x+1) (x-4)

OpenStudy (anonymous):

4 and -1

OpenStudy (irishboy123):

@Adi3 good stuff tip: in an exam, i would actually take those numbers and put them back into the original equation just to be triply sure :p

OpenStudy (adi3):

how would you do that??

OpenStudy (anonymous):

@IrishBoy123 Labhraíonn tú Gaeilge? Ansin, is féidir leat cabhrú liom ... Ní féidir leat?

OpenStudy (adi3):

@IrishBoy123

OpenStudy (adi3):

@IrishBoy123

OpenStudy (adi3):

oh is it foil??

OpenStudy (irishboy123):

check these \(\frac{ -1-1 }{ -1+2} = \frac{ 2 }{ -1 }\) ??? \(\frac{ 4-1 }{ 4+2} = \frac{ 2 }{ 4 }\) ???

OpenStudy (anonymous):

@IrishBoy123 Dhíth orm desperately do chabhair . Níorbh fhéidir leat le do thoil cabhrú liom nuair a fhaigheann tú an deis ? Tá sé éadóchasach .

OpenStudy (adi3):

wht did you just do??

OpenStudy (irishboy123):

plugged your answers into the original equation you were given...

OpenStudy (irishboy123):

ie plug 4 and -1 into \(\frac{ x-1 }{ x+2} = \frac{ 2 }{ x }\) to check them....

OpenStudy (adi3):

how would you plug two equations in x?

OpenStudy (irishboy123):

i will be around and about later, feel free to tag.

OpenStudy (irishboy123):

put them in separately so for x = 4 \(\frac{ x-1 }{ x+2} = \frac{ 2 }{ x } \implies \frac{ 4-1 }{ 4+2} = \frac{ 2 }{ 4 }\) that is \(\frac{3}{6} = \frac{2}{4} \checkmark\)

OpenStudy (adi3):

ok, medal me pls.

OpenStudy (anonymous):

@IrishBoy123 Hilfe bitte?